The area of the rectangle is 80 square centimeters. If the perimeter is 36 centimeters, find the length and width of the rectangle.
Answer by oberobic(2304)

P = 2(L+W)
P = 36
2L + 2W = 36
.
L * W = 80
so
L = 80/W
.
substitute
.
2(80/W) +2W = 36
.
multiply both sides by W to eliminate the fraction
.
2(80) + 2W*W = 36W
.
divide both sides by 2
.
80 + W^2 = 18W
.
W^2 -18W +80 = 0
.
factor
.
(W -8)(W-10) = 0
.
W = 8 or 10
which means
L = 10 or 8
respectively
.
Answer: Length = 10 and Width = 8 OR Length = 8 and Width = 10.
.
Done.